Calculate the surface-area-to-volume ratio of a cube 3 inches on each side.
Answer by mananth(16946)

surface area of a cube = 6x^2 where x is the side
Volume of cube = x^3 where x is the side
surface area = 6*9=54
Volume = 3*3*3= 27
Ratio = 54/27
=2